# Dark Kinetic Heating of Exoplanets and Brown Dwarfs

### v1

@@ V Other Astrophysical Targets
  • White dwarfs: White dwarfs are smaller but more massive than a brown dwarf, and have gravitational capture rates similar to that of a brown dwarf. However, white dwarfs have higher background temperatures, and thus require more dark heating to provide a detectable signal. White dwarfs have previously been considered in the context of dark matter heating Mochkovitch et al. (1985); Moskalenko and Wai (2006); Hooper et al. (2010); McCullough and Fairbairn (2010); Krall and Reece (2018); Curtin and Setford (2021); Panotopoulos and Lopes (2020); Bell et al. (2021); Ramirez-Quezada (2023); DeRocco et al. (2022), but only in Globular clusters where the dark matter content may be high enough to overcome large heat backgrounds (though the dark matter content in considered systems such as Messier 4 is highly uncertain Moore (1996); Saitoh et al. (2006)). Thanks to our dark kinetic heating effect, dark matter signals can be possible in the local position where the dark matter density is instead well known. We consider a benchmark white dwarf with parameters equal to the white dwarf WD 0552-002 Vincent et al. (2023), which has an effective temperature of about 5000 K. Note that the thermal heating equilibrium timescale for white dwarfs was not previously calculated in the literature to our knowledge, and so we detail it in Appendix F.
@@
  • Earth: While the Earth is a much less dense target than the others considered thus far, leading to lower potential capture rates, we are able to robustly measure its heat power. Due to its low escape velocity, the Earth was only considered previously in the context of dark matter annihilation heating Mack et al. (2007); Starkman et al. (1990); Bramante et al. (2020) or asymmetric dark matter that collapses into small evaporating black holes Acevedo et al. (2021). The internal heat power of the Earth has been measured to be $44.2\pm 1$ TW Pollack et al. (1993). As such, we will show the dark energy deposition as a fraction of this value. Note that a long-range dark matter-baryon force can also affect direct detection experiments on Earth, giving complementary signals; this has been previously explored in Refs. Davoudiasl (2017); Davoudiasl et al. (2020).
@@
In Fig. 10, the fact that there are any parameters for which the Earth is most sensitive is a testament both to the strong effect of our signal regardless of target size, and to the single terawatt precision measurements of Earth’s internal power. However, note that at 1 GeV dark matter mass, and the Earth’s maximum capture cross section of about $10^{-33}$ cm2, this may be already ruled out by direct detection, depending on the particle physics model. In Fig. 10 we aim to demonstrate the complementary in the force range and coupling strength for these objects with different properties; at other dark matter masses these comparisons should still qualitatively hold at cross sections not ruled out by direct detection.
@@ Acknowledgements.
We thank Zach Bogorad, Joseph Bramante, Peter Denton, Vincent Lee, Philip Schuster, and Natalia Toro for helpful comments and discussions. JFA, RKL, and AJR are supported by the U.S. Department of Energy under Contract DE-AC02-76SF00515. AJR is also supported by the NSF GRFP under grant DGE-2146755.
Appendix A $\Phi$ Screening in High Dark Matter Densities
@@
Appendix B Maximum Impact Parameter Calculation
@@
For reasons we detail in Appendix C, we will stay within the regime where $|\Phi(r>R)|\lesssim m_{\chi}$. In particular, for $|\Phi|\ll m_{\chi}$, we can drop terms of $\mathcal{O}(|\frac{\Phi}{m_{\chi}}|^{2})$ as well as terms $\propto G\frac{\Phi}{m_{\chi}}$, which are small. Finally, approximating also $\gamma_{\chi}\simeq 1$, since the dark matter particle is non-relativistic far from the celestial body, we obtain Eq. (11) of the main text,
@@
Appendix C Velocity Boost with Scalar Potential
C.1 Velocity Boost Calculation
@@
This indicates that there exists a finite value of $\Phi=-m_{\chi}$ for which the particle seemingly reaches the speed of light. This is of course non-physical, and as $|\Phi|\rightarrow m_{\chi}$, the energy becomes large enough to spontaneously pair-produce, and other radiative effects become significant. We avoid the parameter space where the scalar potential reaches this critical value at the surface of the object throughout its evolution.
@@
C.2 Time Variation of Energy
@@
Appendix D Calculating $\langle{(b_{\rm max}^{\rm dark})^{2}v_{\chi}\rangle}$
@@
Appendix E Initial Dark Matter Growth versus Evaporation
@@
Appendix F White Dwarf Thermal Equilibrium Timescales

### v2

@@ V Other Astrophysical Targets
  • White dwarfs: White dwarfs are smaller but more massive than a brown dwarf, and have gravitational capture rates similar to that of a brown dwarf. However, white dwarfs have higher background temperatures, and thus require more dark heating to provide a detectable signal. White dwarfs have previously been considered in the context of dark matter heating Mochkovitch et al. (1985); Moskalenko and Wai (2006); Hooper et al. (2010); McCullough and Fairbairn (2010); Krall and Reece (2018); Curtin and Setford (2021); Panotopoulos and Lopes (2020); Bell et al. (2021); Ramirez-Quezada (2023); DeRocco et al. (2022), but only in Globular clusters where the dark matter content may be high enough to overcome large heat backgrounds (though the dark matter content in considered systems such as Messier 4 is highly uncertain Moore (1996); Saitoh et al. (2006)). Other recent analyses have explored the ignition of Type-Ia supernovae from localized heating by heavy asymmetric point-like dark matter Bramante (2015); Acevedo and Bramante (2019); Janish et al. (2019) and composite states Graham et al. (2018); Acevedo et al. (2021a); Raj (2024). Thanks to our dark kinetic heating effect, dark matter signals can be possible in the local position where the dark matter density is instead well known. We consider a benchmark white dwarf with parameters equal to the white dwarf WD 0552-002 Vincent et al. (2023), which has an effective temperature of about 5000 K. Note that the thermal heating equilibrium timescale for white dwarfs was not previously calculated in the literature to our knowledge, and so we detail it in Appendix G.
@@
  • Earth: While the Earth is a much less dense target than the others considered thus far, leading to lower potential capture rates, we are able to robustly measure its heat power. Due to its low escape velocity, the Earth was only considered previously in the context of dark matter annihilation heating Mack et al. (2007); Starkman et al. (1990); Bramante et al. (2020) or asymmetric dark matter that collapses into small evaporating black holes Acevedo et al. (2021b). The internal heat power of the Earth has been measured to be $44.2\pm 1$ TW Pollack et al. (1993). As such, we will show the dark energy deposition as a fraction of this value. Note that a long-range dark matter-baryon force can also affect direct detection experiments on Earth, giving complementary signals; this has been previously explored in Refs. Davoudiasl (2017); Davoudiasl et al. (2020).
@@
In Fig. 10, the fact that there are any parameters for which the Earth is most sensitive is a testament both to the strong effect of our signal regardless of target size, and to the single terawatt precision measurements of Earth’s internal power. However, note that at 1 GeV dark matter mass, and the Earth’s maximum capture cross section of about $10^{-33}$ cm2, this may be already ruled out by direct detection, depending on the particle physics model. In Fig. 10 we aim to demonstrate the complementarity in the force range and coupling strength for these objects with different properties; at other dark matter masses these comparisons should still qualitatively hold at cross sections not ruled out by direct detection.
@@ Acknowledgements.
We thank Zach Bogorad, Joseph Bramante, Peter Denton, Vincent Lee, Nirmal Raj, Philip Schuster, Juri Smirnov, and Natalia Toro for helpful comments and discussions. JFA, RKL, and AJR are supported by the U.S. Department of Energy under Contract DE-AC02-76SF00515. AJR is also supported by the NSF GRFP under grant DGE-2146755.
Appendix A Self-Interaction Jeans Length
We show that the self-interaction considered does not lead to Jeans instabilities on scales of order the interaction range $\lambda$ for halo dark matter. We begin by estimating the free-fall time of a dark matter sphere under this new force, which is given by
$t_{\rm ff}\simeq\sqrt{\frac{3\pi m_{\chi}^{2}}{32\alpha_{\chi}\rho_{\chi}}}\simeq 1.3\times 10^{9}\ {\rm s}\,\left(\frac{\alpha_{\chi}}{10^{-26}}\right)^{-1/2}\left(\frac{\rho_{\chi}}{\rm GeV\ cm^{-3}}\right)^{-1/2}\left(\frac{m_{\chi}}{\rm GeV}\right)\leavevmode\nobreak\ .$ (35)
This timescale must be compared to the sound crossing time over the same region; if the sound waves take longer than $t_{\rm ff}$ to cross a region of size $\lambda$, they will be unable to stabilize the clump against collapse via the attractive dark force. Given the extremely weak couplings we analyze, the sound speed of the dark matter $c_{s}$ can be approximated as that of an ideal gas
$c_{s}\simeq v_{p}\simeq 270\ \rm km\ s^{-1}\,,$ (36)
and, therefore, the sound crossing time across a region of size $\lambda$ is approximately
$\left(\frac{c_{s}}{\lambda}\right)^{-1}\simeq 2.5\times 10^{5}\ {\rm s}\,\left(\frac{\lambda}{100R_{\odot}}\right)\left(\frac{v_{p}}{270\ \rm km\ s^{-1}}\right)^{-1}\,,$ (37)
which is always much shorter than the free-fall time for $\lambda\lesssim 100\,R_{\odot}$. Therefore, across the range of couplings, dark matter masses, and mediator ranges considered, we do not expect the new dark force to induce a Jeans instability in the Galactic halo dark matter. More broadly, new effects other than dark kinetic heating may be observable, and would be interesting to consider as new dark sector discovery avenues, but are outside the scope of this work.
Appendix B $\Phi$ Screening in High Dark Matter Densities
@@
Appendix C Maximum Impact Parameter Calculation
@@
For reasons we detail in Appendix D, we will stay within the regime where $|\Phi(r>R)|\lesssim m_{\chi}$. In particular, for $|\Phi|\ll m_{\chi}$, we can drop terms of $\mathcal{O}(|\frac{\Phi}{m_{\chi}}|^{2})$ as well as terms $\propto G\frac{\Phi}{m_{\chi}}$, which are small. Finally, approximating also $\gamma_{\chi}\simeq 1$, since the dark matter particle is non-relativistic far from the celestial body, we obtain Eq. (11) of the main text,
@@
Appendix D Velocity Boost with Scalar Potential
D.1 Velocity Boost Calculation
@@
This indicates that there exists a finite value of $\Phi=-m_{\chi}$ for which the particle seemingly reaches the speed of light. This is of course non-physical and as $|\Phi|\rightarrow m_{\chi}$, other physical effects become significant that are outside the scope of this work. We avoid the parameter space where the scalar potential reaches this critical value at the surface of the object throughout its evolution.
@@
D.2 Time Variation of Energy
@@
Appendix E Calculating $\langle{(b_{\rm max}^{\rm dark})^{2}v_{\chi}\rangle}$
@@
Appendix F Initial Dark Matter Growth versus Evaporation
@@
Appendix G White Dwarf Thermal Equilibrium Timescales